### v3

@@ 5.4 Results during the Covid-19 Pandemic
The results presented so far gave an overall picture of the accuracy of the various models across the different forecast horizons, focusing on the evaluation sample January 1985 to December 2019. In this section, we focus on the post Covid-19 period sample, Jan 2020 to December 2022 (to reduce the impact on the results of the extreme outliers right around the onset of the pandemic, we excluded from our analysis the forecasts made between March and June 2020). [^16] The choice of excluding from the sample the most problematic Covid-19 related observation is based exclusively on simplicity. Still, there is evidence that simple approaches such as this might work well, see Schorfheide and Song (2021). One could also consider more sophisticated ways to deal with the Covid-19 observations such as Carriero et al. (2022) or Lenza and Primiceri (2022).
Figure 4 and Figure 5 replicate our pre-Covid-19 analysis, again first comparing the forecast accuracy of the various TSLMs, and next stacking that against the BVARs and factor model. Compared to what we shown in Figure 2 and Figure 3 there is an overall deterioration in forecast accuracy for all the models, which all struggle in overperforming the benchmark. Within the LLMs, only Moirai-large stands out, with a performance comparable with that of the factor model and the BVAR with symmetric conjugate priors. It’s worth pointing out that in this sample the econometric models tend to make at times large forecast errors, mainly during the second half of 2020. In the case of the factor model it happens especially for interest rates and spreads.
@@
The usual caveat about TSLMs should be kept in mind, i.e. the fact that, as shown in Table 1, these models do include the pandemic period in their training sample (and a few of these models also include may of the macroeconomic variables we set out to forecasts), and this makers it harder to pin down to which extent this forecasting performance could have been achieved in real time.
@@ 5.5 How do TSLMs Work with Persistent Series?
As described above, some of the series are transformed before entering the estimation stage of the models. Looking at the results shown so far, there appears to be some indication that TSLMs perform relatively better when the degree of persistence of the series they utilize is relatively low. To further investigate this aspect, we break down the previous results by separately looking at series with different degrees of persistence. Specifically, Figure 6 shows results for variables for which the one lag partial autocorrelation is below 0.9, while Figure 7 shows results for variables for which the one lag partial autocorrelation is above 0.9. The latter variables are of course highly persistent.
@@
Turning the attention to TSLMs, for series with low to moderate persistence, they seem to perform relatively better, even though it should be noted that they are still outperformed by factor models for this subset of series. The overall picture changes starkly when considering highly persistent variables. In these cases, one can observe a definite increase in the variation of the performance of TSLM.
@@
To give an example of the extent of variation in performance, Figure 8 shows the time series of out-of-sample forecasts produced by the factor model, the BVAR (with symmetric priors) and TimesFM model for a few selected variables. While for the 1-step ahead horizons the forecasts are quite similar (left side panels), large differences emerge at the 12-step ahead horizon (right side panels).
@@
Figure 8: Time series of out-of-sample forecasts produced by the factor model, the BVAR with symmetric priors and the timesFM model, for selected variables. Left panels show 1-step ahead forecasts, and right panels show 12-step ahead forecasts. The sample is January 1985 to December 2019.
@@
For Housing Starts (HOUST), the 12-step ahead forecasts produced by TimesFM are remarkably good, as this model is able to almost perfectly predict the pronounced slowdown that happened during the 2007 financial crisis and the ensuing recession. This results in an astonishing 50 percent improvement in RMSFE compared to the AR benchmark. Instead, the econometric models only pick up this slowdown with a delay, and the factor model even fails to recognize the depth of the trough. Of course, one must keep in mind that TimesFM includes this recessionary time period in its training sample, so it seems to be on some level aware that a change in time series behavior is to be expected within these dates. For weekly hours (CES0600000007) TimesFM outperforms the factor model but underperforms the BVAR. For variables such as the spread between the 10 year interest rate and the Federal Funds rate (T10YFFM) and the spread between corporate bonds and the Federal Funds rate (BAAFFM), the performance of TSLMs is subpar, featuring entire periods of "hallucination" in the second half of the 80s and in 2017-2019, resulting in RMSFE losses of the order of 15-20 percent.
@@
6 Fine-tuning
Here, we selected the two best models from our zero-shot experiments – TimesFM and Moirai-large – and fine-tuned them on FRED-MD data. Note that some of the pretrained TSLMs, including Moirai, are trained on datasets that overlap with the FRED-MD. However, the influence of any single dataset in the pretraining phase gets diluted within the vast and diverse pretraining corpus. Consequently, the models learn general time series patterns, rather than specializing in the specific characteristics of FRED-MD. To make a TSLM more dataset-specific, fine-tuning is required. The main challenge we face in this context is that FRED-MD is a relatively small dataset, which may lead to overfitting problems. [^18] Kumar et al. (2022) discusses how in these situations the model may also lose the knowledge acquired during pretraining, leading to the so-called catastrophic forgetting. To address these challenges, we employ a simple parameter-efficient fine-tuning strategy as opposed to continued-pretraining. In particular, we freeze the transformer layers of the two TSLMs, thereby reducing the number of trainable parameters. [^19] Kumar et al. (2022) refers to this as Linear Probing. Note that freezing the pretrained model’s weights (parameters) is a special case of LoRA, which additionally injects trainable rank decomposition matrices into each layer of TSLM. After freezing the transformer layers, the number of trainable parameters reduced to approximately $\leq 10$M for both TimesFM (instead of $200$M) and Moirai-large (instead of $311$M), primarily part of the final layers. [^20] In the case of Moirai, we also consider fine-tuning using the continued-pretraining strategy, as this was the recommended approach provided by the authors in the model’s repository, but in this case the model ended up severely overfitting and producing poor out-of-sample forecasts. Throughout, to keep computational costs manageable, we chose to fine-tune each model once every 12 months, and the resulting models were then used for zero-shot forecasting over the subsequent 12 months. [^21] Within each fine-tuning round, the training FRED-MD data was split into an 80:20 ratio, with 20% reserved for validation. We trained these models for 200 epochs, utilizing early stopping with a patience of 5 to mitigate overfitting. To further counter overfitting and ensure stable training on our limited dataset, we reduced the learning rate, preventing large weight updates. We run the fine-tuning experiment on a stock server equipped with a single NVIDIA A6000 GPU. In terms of wall-clock time, fine-tuning took about 10 hours for each of the two models. Hence, comparatively, fine-tuning remains significantly more resource-intensive than the BVARs and factor models we considered, which can be trained on standard CPUs in a fraction of the time. [^22] We also considered the continued-pretraining option for the same models, and in that case the total computing time was approximately 90 hours.
@@ 7 Conclusions
Despite these limitations, our results indicate that TSLMs exhibit some promising features, particularly in capturing nonlinearities and adapting to evolving economic conditions. Their relatively stronger performance in the post-Covid-19 period suggests that they may have advantages in handling structural breaks and rapidly changing economic environments. As research in this area continues to evolve, it may be worthwhile exploring hybrid approaches that integrate the flexibility and generality of foundational models with the theoretical rigor of econometric methods. For example, it would be interesting to investigate whether incorporating structured economic constraints into the TSLMs’ architectures could enhance their reliability. We leave this topic to further research.
@@ A.1 Prior Moments of Natural Conjugate Prior
where $\Phi_{l}^{(ij)}$denotes the element in position $(i,j)$ in the matrix $\Phi_{l}$. The prior mean $\Phi^{\ast}$ is set to either $1$ or 00 depending on the presumed order of integration of the series. For the intercept ($l=c$) we assume an uninformative prior with mean 00 and standard deviation $\lambda_{0}\sigma_{i}$. The shrinkage parameter $\lambda_{1}$ measures the overall tightness of the prior: when $\lambda_{1}\rightarrow 0$ the prior is imposed exactly and the data do not influence the estimates, while as $\lambda_{1}\rightarrow\infty$the prior becomes loose and the prior information does not influence the estimates, which will approach the standard $OLS$ estimates. To set each scale parameter $\hat{\sigma}_{j}$ in (A.1) we follow common practice (see e.g. Litterman (1986); Sims and Zha (1998)) and set it equal to the standard error of regression from a univariate autoregressive model. The parameters $\sigma_{i}^{2}$ appearing in (A.1) are instead coming directly from the error variance $\Sigma$ on which this prior is conditioning upon, trough the Kronecker multiplication shown in (4). [^23] Incidentally, note that such multiplication also induces correlation across coefficients belonging to different equations, which is not apparent in (A.1) because it focuses only on the variances of the individual coefficients.

### v4

@@ 5.3 Results during the Covid-19 Pandemic
The results presented so far gave an overall picture of the accuracy of the various models across the different forecast horizons, focusing on the evaluation sample January 1985 to December 2019. In this section, we focus on the post Covid-19 period sample, Jan 2020 to December 2022. Also, we exclude from our analysis the forecasts made between March and June 2020, to reduce the impact on the results of the extreme outliers right around the onset of the pandemic. [^19] The choice of excluding from the sample the most problematic Covid-19 related observation is based exclusively on simplicity. Still, there is evidence that simple approaches such as this might work well, see Schorfheide and Song (2021). One could also consider more sophisticated ways to deal with the Covid-19 observations such as Carriero et al. (2022) or Lenza and Primiceri (2022).
Figure 4 compares the forecast accuracy of the various TSLMs, and Figure 5 considers the best performing TSLMs against the various econometric models.
Compared to what we shown in Figure 2 and Figure 3 there is an overall deterioration in forecast accuracy for all the models, which all struggle in overperforming the benchmark. Within the LLMs, only Moirai-large stands out, with a performance comparable with that of the factor model and the BVAR. In this sample the econometric models tend to make at times large forecast errors, mainly during the second half of 2020. For the factor model it happens especially for interest rates and spreads.
The usual caveat applies that, as shown in Table 1, namely that the TSLMs include the pandemic period in their training sample (and a few also include many of the variables we set out to forecasts), and this makes it harder to pin down to which extent this forecasting performance could have been achieved in real time.
6 Understanding the Drivers of TSLM Forecast Performance
To better understand the results presented so far, this Section explores three key aspects of the TSLMs and how these impact the models’ forecast performance. In Section 6.1 we consider the effects of possible "data leakage", which as we mentioned before can spuriously tilt the forecast accuracy scale in favor of the TSLMs. In Section 6.2 we show that highly persistent series are particularly problematic for TSLMs to forecast. Finally, in Section 10 we consider the impact of fine tuning on the TSLM forecasts, i.e. updating the model parameter estimates to reflect more closely the information and data patterns of the macroeconomic time series contained in the FRED-MD database.
@@ 6 Understanding the Drivers of TSLM Forecast Performance
6.1 Investigating the Impact of Different Information Sets on Forecast Accuracy
Figure 6: Distribution of RMSFEs (relative to an AR benchmark) for the econometric models and a selection of TSLMs. BVARs and Factor models are trained using data up to December 2022. The evaluation sample is January 1985 to December 2019.
In the Introduction, we emphasized that it is not straightforward to assess the real-time forecast accuracy of TSLMs, mainly because these models are trained on data lacking timestamps. Thus, previous results may overstate TSLM forecasting ability, introducing a potential look-ahead bias due to some of the training data postdating the forecast date. To further complicate things, we saw in Table 1 that some of the TSLM models do include in the training set the FRED-MD database, which means they potentially have access to the perfect forecast. Although retraining the TSLM models in real time is difficult and costly, we can perform the opposite experiment of giving the same unfair advantage to the econometric models. This is the experiment we describe in this subsection.
Specifically, we re-estimated the econometric models using the entire sample available, up to December 2022. This date precedes or coincides with the dates in which the TSLMs were trained (based on the year of publication of the corresponding papers, during 2024). While this comparison is not perfect, it can provide a first order approximation of the potential impact of using different information sets in the model training phase. Figure 6 shows the results of this experiment, where the BVAR and Factor model RMSFEs are replaced by their full-sample counterparts. Moirai and TimesFM results, taken from Figure 2, are included to make the comparison easier.
The results are revealing. The forecasting performance of BVAR and Factor models is dramatically enhanced by the use of the extra information, with their respective box plots lying entirely below one. In other words, these models consistently beat the AR benchmark across most series and horizons. They also perform significantly better than the TSLMs, which appeared to perform comparably to the econometric models only when the latter were estimated *without* the benefit of hindsight.
@@ 6.2 How do TSLMs Work with Persistent Series?
As described above, many of the macroeconomic series are transformed before entering the estimation stage of the models. Looking at the results shown so far, there appears to be some indication that TSLMs perform relatively better when the degree of persistence of the series they utilize is relatively low. To further investigate this aspect, we break down the previous results by separately looking at series with different degrees of persistence. Specifically, Figure 7 shows results for variables for which the one lag partial autocorrelation is below 0.7, while Figure 8 shows results for variables for which the one lag partial autocorrelation is above 0.7. The latter variables are of course highly persistent.
@@
Turning the attention to TSLMs, for series with low to moderate persistence, they seem to perform relatively better, still on par with BVARs and factor models. The overall picture changes starkly when considering highly persistent variables. In these cases, one can observe a definite increase in the variation of the performance of TSLM. Interestingly, the same pattern applies to the non-linear econometric models, NNAR and BART.
@@
To give an example of the extent of variation in performance, Figure 9 shows the time series of out-of-sample forecasts produced by the factor model, the BVAR and TimesFM for a few selected variables. While for the 1-step ahead horizon the forecasts are quite similar (left side panels), large differences emerge at the 12-step ahead horizon (right side panels).
@@
Figure 9: Time series of out-of-sample forecasts produced by the factor model, the BVAR, and the TimesFM model for selected variables. Left panels show 1-step ahead forecasts, and right panels show 12-step ahead forecasts. The sample is January 1985 to December 2019.
@@
For Housing Starts (HOUST), the 12-step ahead forecasts produced by TimesFM are remarkably good, as this model is able to almost perfectly predict the pronounced slowdown that happened during the 2007 financial crisis and the ensuing recession. This results in an astonishing 50 percent improvement in RMSFE compared to the AR benchmark. Instead, the econometric models only pick up this slowdown with a delay, and the factor model even fails to recognize the depth of the trough. Of course, one must keep in mind that TimesFM includes this recessionary time period in its training sample, so it seems to be on some level aware that a change in time series behavior is to be expected within these dates. For weekly hours (CES0600000007) TimesFM outperforms the factor model but underperforms the BVAR. For variables such as the spread between the 10 year interest rate and the Federal Funds rate (T10YFFM) and the spread between corporate bonds and the Federal Funds rate (BAAFFM), the performance of TSLMs is subpar, featuring entire periods of "hallucination" in the second half of the 80s and in 2017-2019, resulting in RMSFE losses on the order of 15-20 percent.
6.3 Fine-tuning
Here, we selected the two best models from our zero-shot experiments – TimesFM and Moirai-large – and fine-tuned them on FRED-MD data. Note that some of the pretrained TSLMs, including Moirai, are trained on datasets that overlap with the FRED-MD. However, the influence of any single dataset in the pretraining phase gets diluted within the vast and diverse pretraining corpus. Consequently, the models learn general time series patterns, rather than specializing in the specific characteristics of FRED-MD. To make a TSLM more dataset-specific, fine-tuning is required. The main challenge we face in this context is that FRED-MD is a relatively small dataset, which may lead to overfitting problems. [^21] Kumar et al. (2022) discusses how in these situations the model may also lose the knowledge acquired during pretraining, leading to the so-called catastrophic forgetting. To address these challenges, we employ a simple parameter-efficient fine-tuning strategy as opposed to continued-pretraining. In particular, we freeze the transformer layers of the two TSLMs, thereby reducing the number of trainable parameters. [^22] Kumar et al. (2022) refers to this as Linear Probing. Note that freezing the pretrained model’s weights (parameters) is a special case of LoRA, which additionally injects trainable rank decomposition matrices into each layer of TSLM. After freezing the transformer layers, the number of trainable parameters reduced to approximately $\leq 10$M for both TimesFM (instead of 200200M) and Moirai-large (instead of 311311M), primarily part of the final layers. [^23] In the case of Moirai, we also consider fine-tuning using the continued-pretraining strategy, as this was the recommended approach provided by the authors in the model’s repository, but in this case the model ended up severely overfitting and producing poor out-of-sample forecasts. Throughout, to keep computational costs manageable, we chose to fine-tune each model once every 12 months, and the resulting models were then used for zero-shot forecasting over the subsequent 12 months. [^24] Within each fine-tuning round, the training FRED-MD data was split into an 80:20 ratio, with 20% reserved for validation. We trained these models for 200 epochs, utilizing early stopping with a patience of 5 to mitigate overfitting. To further counter overfitting and ensure stable training on our limited dataset, we reduced the learning rate, preventing large weight updates. We run the fine-tuning experiment on a stock server equipped with a single NVIDIA A6000 GPU. In terms of wall-clock time, fine-tuning took about 10 hours for each of the two models. Hence, comparatively, fine-tuning remains significantly more resource-intensive than the BVARs and factor models we considered, which can be trained on standard CPUs in a fraction of the time. [^25] We also considered the continued-pretraining option for the same models, and in that case the total computing time was approximately 90 hours.
@@ 6.3 Fine-tuning
7 A TSLM-Augmented BVAR (TSLM-BVAR)
The results so far show that although the econometric model consistently outperforms the AR benchmark, in some cases TSLMs can be very accurate. At the same time, in a number of cases the performance of TSLMs can be significantly sub-par. One way to rationalize these results is that while TSLMs are capable of exploiting vast datasets, they still lack the distilled domain knowledge built from decades of macroeconomic analysis. Still, in some cases, their ability to capture nonlinearities in the data can be quite advantageous.
In this section, we consider an approach that can help leverage the best of both worlds: that is, exploiting the domain knowledge and simplicity of the econometric models while relying on TSLMs to capture nonlinear behavior missed by the traditional approaches.
Our proposal is to build a simple hybrid approach. First, an econometric model (we use the BVAR) is fit to the data and the in-sample residuals are computed. Second, the residuals from the first step are modeled using a TSLM. Finally, forecasts are computed from both models and combined. Hence, the final forecast adjusts the econometric prediction by adding the predicted residual component, effectively correcting systematic biases and capturing nonlinearities.
Formally, let $y_{i,t}$ denote the target variable of interest. The procedure comprises the following steps:
  1. 1.
    Econometric Model Forecast. Estimate a BVAR using historical data on all series and generate forecasts:
    $\hat{y}_{i,t+h}^{(E)}=\mathbb{E}\bigl[y_{i,t+h}\,\big|\,\mathcal{I}_{t}\bigr],$ (12)
    where $\mathcal{I}_{t}$ denotes the information set up to time $t$. Compute the in-sample residuals:
    $\hat{e}_{i,\tau}=y_{i,\tau}-\hat{y}_{i,\tau}^{(E)},\quad\tau=1,...,t.$ (13)
  2. 2.
    Foundational Model Training and forecasting. Train a foundational time series model on the residuals $\{\widehat{e}_{i,\tau}\}_{\tau=1}^{t}$ to learn remaining patterns and then generate a zero-shot forecast:
    $\hat{e}_{i,t+h}^{(F)}=\mathbb{E}\bigl[\hat{e}_{i,t+h}\,\big|\,\mathcal{H}_{t}\bigr],$ (14)
    where $\mathcal{H}_{t}$ denotes the residual history up to time $t$.
  3. 3.
    Hybrid Forecast Construction. The final hybrid forecast combines the econometric prediction and the TSLM forecast:
    $\hat{y}_{i,t+h}^{(\text{H})}=\hat{y}_{i,t+h}^{(E)}+\hat{e}_{i,t+h}^{(F)}.$ (15)
This approach retains the interpretability of the BVAR while flexibly capturing nonlinear residual dynamics through the foundational architecture.
Figure 11 and Table 4 present the results of this analysis. In particular, Figure 11 shows results for the original BVAR plus the two best-performing TSLMs, Moirai and TimesFM, and finally two hybrid versions combining the BVAR with both TSLMs. The hybrid approach shifts the TSLM error distribution downward, placing most of its mass below one—the threshold for outperforming the AR benchmark.
Figure 11: Distribution of RMSFEs (relative to an AR benchmark) for the original BVAR and the new procedure combining the BVAR with an error correction based on the two best-performing TSLMs, TimesFM and Moirai-large. The evaluation sample is January 1985 to December 2019.
h=1 h=3
Median Std Min Max Median Std Min Max
BVAR 0.983 0.060 0.827 1.158 0.970 0.060 0.767 1.089
Moirai-large 0.999 0.102 0.703 1.436 0.978 0.099 0.637 1.158
Moirai-large hybrid 0.973 0.075 0.734 1.139 0.960 0.079 0.674 1.095
TimesFM 1.014 0.129 0.706 1.482 0.980 0.127 0.635 1.318
TimesFM hybrid 0.980 0.081 0.720 1.150 0.956 0.089 0.670 1.105
h=6 h=12
Median Std Min Max Median Std Min Max
BVAR 0.991 0.068 0.725 1.106 0.999 0.080 0.671 1.162
Moirai-large 0.991 0.096 0.600 1.159 1.001 0.113 0.619 1.324
Moirai-large hybrid 0.983 0.079 0.648 1.102 0.998 0.087 0.671 1.164
TimesFM 0.990 0.142 0.593 1.629 1.001 0.158 0.482 1.440
TimesFM hybrid 0.984 0.080 0.636 1.119 0.999 0.088 0.677 1.174
Table 4: Summary statistics of RMSFE ratios by model and forecast horizon, including median, standard deviation, minimum, and maximum values.
Moreover, Table 4 shows that the maximum relative RMSFEs from the TSLMs are greatly reduced by the hybrid approach: at one step ahead, the error of Moirai decreases from 1.436 to 1.139, and that of TimesFM from 1.482 to 1.150. These figures are very close to the maximum error of the BVAR, which is 1.158. Thus, the use of econometric models in the first step substantially reduces the risk of hallucinations and other aberrant forecast behaviors. Incorporating the domain knowledge of econometric models helps TSLMs across the board by lowering forecast errors for all series.
When comparing the hybrid versions with the BVAR (first box on the left in the figure), one can notice that adding to this model the predictions of the residuals via TSLMs does not adversely affect the right tail of the distribution, as there are no forecasts that worsen as a result. Instead, the addition has a positive effect on the left tail: for those forecasts that were already improving over the AR benchmark, the magnitude of the improvement increases. In Table 4, the best relative RMSFE from the BVAR model shows improvements ranging from 5% to 12% depending on the forecast horizon. For example, at 1-step ahead the best relative RMSFE drops from 0.827 to 0.734 or 0.720 when the BVAR is used in conjunction with Moirai and TimesFM, respectively.
Overall, the hybrid approach improves upon both the econometric model and the TSLMs used in isolation. Using the econometric model in the first step prevents TSLMs from producing subpar forecasts, while using the TSLMs on the residuals enhances the performance of the econometric model. This suggests that the domain knowledge of econometric models is robust and reliable, but can be further improved using TSLMs to capture nonlinearities. [^26] While these results are encouraging, one must not forget the caveat that the performance improvement form TSLMs we observed here might be due to the fact it has a spuriously superior information set, as discussed in Section 6.1.
@@ 8 Conclusions
Despite these limitations, our results indicate that TSLMs exhibit some promising features, particularly in capturing nonlinearities and adapting to evolving economic conditions. Their relatively stronger performance in the post-Covid-19 period suggests that they may have advantages in handling structural breaks and rapidly changing economic environments.
With this in mind, we proposed a TSLM-augmented BVAR (TSLM-BVAR) hybrid approach leverages the complementary strengths of econometric models and foundational time series learning. By first generating forecasts with a BVAR and then modeling the residuals using TSLMs, the method captures nonlinear patterns missed by traditional models while retaining the interpretability and domain knowledge of econometrics. Empirical results show that this hybrid procedure consistently reduces forecast errors compared to either component in isolation, particularly mitigating extreme errors in TSLM predictions. Overall, the TSLM-BVAR approach demonstrates that integrating structured econometric information with flexible machine learning models can substantially improve forecasting performance.
@@ A.1 Prior Moments of Natural Conjugate Prior
where $\Phi_{l}^{(ij)}$denotes the element in position $(i,j)$ in the matrix $\Phi_{l}$. The prior mean $\Phi^{\ast}$ is set to either $1$ or 0 depending on the presumed order of integration of the series. For the intercept ($l=c$) we assume an uninformative prior with mean 0 and standard deviation $\lambda_{0}\sigma_{i}$. The shrinkage parameter $\lambda_{1}$ measures the overall tightness of the prior: when $\lambda_{1}\rightarrow 0$ the prior is imposed exactly and the data do not influence the estimates, while as $\lambda_{1}\rightarrow\infty\$the prior becomes loose and the prior information does not influence the estimates, which will approach the standard $OLS$ estimates. To set each scale parameter $\hat{\sigma}_{j}$ in (A.1) we follow common practice (see e.g. Litterman (1986); Sims and Zha (1998)) and set it equal to the standard error of regression from a univariate autoregressive model. The parameters $\sigma_{i}^{2}$ appearing in (A.1) are instead coming directly from the error variance $\Sigma$ on which this prior is conditioning upon, trough the Kronecker multiplication shown in (4). [^27] Incidentally, note that such multiplication also induces correlation across coefficients belonging to different equations, which is not apparent in (A.1) because it focuses only on the variances of the individual coefficients.